Archer Company is a wholesaler of custom-built air-conditioning units for commercial buildings. It gathered the following monthly data relating to units shipped and total shipping expense:

Month Units Shipped Total Shipping Expense
January 2 $ 2,200
February 5 $ 2,700
March 3 $ 2,100
April 4 $ 2,400
May 6 $ 2,700
June 7 $ 3,700
July 1 $ 1,600

Using the high-low method, estimate the cost formula for shipping expense.

Variable shipping expense: (per unit)

Fixed shipping expense: (per month)
Answer #1

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (3,700-1,600)/(7-1)

= 2,100/6

= $350 per unit

Fixed cost = Highest activity cost - Highest activity x Variable cost per unit

= 3,700 - 7 x 350

= 3,700-2,450

= $1,250 per month

Variable shipping expense: $350 (per unit)

Fixed shipping expense: $1,250 (per month)

Kindly comment if you need further assistance. Thanks‼!